OpenStudy (anonymous):

Simplify the product using the distributive property. (-4h +2)(3h +7) A. -12h2 - 22h + 14 B. -12h2 - 34h - 14 C. -12h2 + 34h - 14 D. -12h2 + 22h + 14

hero (hero):

\(a(b + c) = ab + ac\) \((-4h + 2)(3h + 7) = -4h(3h + 7) + 2(3h + 7)\) \(=-12h^2 -28h + 6h + 14\) \(=-12h^2 -22h + 14\)